# Inhomogeneous mixing: From microscopic dynamics to mesoscopic staircases

### v1

@@ 4.1 Models on the $E\times B$ staircase
Many theoretical models have been suggested to understand the physics of the $E\times B$ staircase. The $E\times B$ staircase was characterized as a mesoscopic transport state with co-existing super-diffusive transport events (avalanches) and permeable transport barriers (zonal flow layers) [67]. It was found in different frameworks [93, 94, 95] that such a state can be realized, though their physical interpretations may differ. In Garbet et al. [93], avalanches and zonal flow layers are the radially propagating wave packets and the wave trapping [96, 97, 98] quasi-periodic zonal flow solution in the system, respectively. In Milovanov et al. [94], zonal flow layers were described by coupled oscillators of nonlinear Schrödinger equation whose vibrations and sub-quadratic interaction allow a super-diffusive radial transport. In Chen et al. [95], drift wave solitons (akin to avalanches) can develop via the nonlinear drift wave and zonal flow interaction and they are confined by spontaneously excited zonal flows.
@@
Some researches [85, 99, 86, 100, 68, 101, 102] have emphasized a role of mixing processes in the formation of mesoscopic zonal flows or profile corrugation. In Kosuga et al. [85], the pressure corrugation was understood as the jamming of avalanching heat flux due to the finite time delay [99]. In Ashourvan and Diamond [86] and Guo et al. [100], the bistability by dynamic mixing length was adopted to implement the inhomogeneous mixing and result in the density corrugation. In Wang et al. [68], the initial zonal flow driven by the Reynolds stress was fortified by the in-phase mean $E_{r}$ variation during the temperature relaxation event. In Leconte and Kobayashi [101], the radial variation of the transport cross phase between the density and electric potential fluctuations (and the associated particle flux) can result in the density corrugation. In Yan and Diamond [102], alternating the enhanced and mitigated transport regions via the wave-particle resonance and non-resonance, respectively, due to the spatial structure of the mean flow lead to a corrugated profile in the trapped ion mode turbulence.
@@ 4.2 Some lessons from analogies
Revisiting details of specific models and comparing them with experimental results are left for future work. Here, in line with the theme of the workshop, “Layering—A structure formation mechanism in oceans, atmospheres, active fluids and plasmas”, the $E\times B$ staircase is viewed as one of self-organized layering phenomena in non-equilibrium complex systems and compared with other long-studied phenomena. Although there may be no universal laws for self-organized structures in complex systems, lessons can be learned by comparing observations in different systems [105].
@@
The $E\times B$ staircase is a pattern in the near-marginal regime characterized by the Kubo number ($Ku=\tau_{\mathrm{corr}}/\tau_{\mathrm{turn}}$, the ratio between the correlation time and the eddy turnover time) larger than 1. $Ku>1$ means more efficient eddy-particle interaction than decorrelation or detrapping processes, and results in the discrepancy from the local diffusive transport models [71]. Both the $E\times B$ staircase and the RB convection cells are patterns that form away from a state dictated by the homogeneous mixing of diffusive process (see Table 4). It demands us to go beyond the diffusive paradigm [36].
However, the $E\times B$ staircase is the secondary structure out of turbulent fluctuations, while the RB cells are more like steady laminar flows in the pattern of regular cells driven by the primary instability. While the scale of the RB convection cells is largely determined by the most unstable linear mode in the system, the nonlinear dynamics would matter for the scales of the $E\times B$ staircase, $\delta^{\mathrm{flow}}$ and $\Delta$. For example, $\delta^{\mathrm{flow}}$ can be extended by the turbulence spreading [107].
As the $E\times B$ staircase can be destroyed by avalanches (turbulent heat flux), the RB cells also become more complex, deformed or destructed, with turbulent fluctuations as the temperature difference is further increased (or, equivalently the viscosity is reduced). In Ramirez and Diamond [108], the effect of turbulent fluctuation on a staircase state of cellular flows was investigated, providing insight into the fluctuation strength condition that the staircase state can survive.
@@
Rayleigh-Bénard cells Zonal jets The $E\times B$ staircase
@@
Key condition $Ra=\frac{\tau_{\mathrm{cond}}}{\tau_{\mathrm{conv}}}>Ra_{\mathrm{crit}}$ $\tau_{\mathrm{wave}}\sim\tau_{\mathrm{turn}}$ $Ku=\frac{\tau_{\mathrm{corr}}}{\tau_{\mathrm{turn}}}>1$
@@
A more relevant and famous analogy with zonal jets in the geophysical fluid has been made [62, 109]. Indeed, the formal analogy can be made between the Charney equation for the (beta-plane) geophysical fluid dynamics and the Hasegawa-Mima equation for the (slab geometry) drift wave turbulence in plasmas.
In those simplified systems, there exists a conserved quantity called potential vorticity (PV; $Q$) attached to a fluid element, i.e. $\frac{dQ}{dt}\approx 0$. For the total PV variation is preserved within the system, if the PV mixing (by turbulent eddies or wave breaking) occurred in one location and reduce the PV gradient there, the PV gradient should be increased in other place. This is called the inhomogeneous PV mixing or PV staircase [65].
@@
One enlightenment from the analogy with the zonal jets is the non-acceleration theorem [111] that implies the important role of turbulence spreading flux (see Section 2.2.2) in accelerating the zonal flow [87, 88]. For example, the turbulence influx through the separatrix can play a role to drive zonal flows in the core plasma (see Section 2.2.4). Turbulence spreading can mediate the interaction between zonal flow layers [89] of the $E\times B$ staircase, suggesting an explanation for its long range order.
In realistic fusion plasmas, however, directly adopting the conventional picture of the inverse cascade is debatable. It relies on the large separation between the injection scale and the dissipation scale, but these scales may not be clearly distinguished in fusion plasmas with multi-scale instabilities. Also, there can be additional processes generating a radial mixing zone in fusion plasmas [112, 73, 74], and defining the corresponding scale to the Rhines scale is not trivial. Avalanches are intrinsically the scale-free transport [70] unless they are limited by zonal flows. The width of the radial mixing region $\Delta$, the radial transport scale or the distance between the zonal flow layers, might be rather set by the competition or interaction between the radial (avalanches) and poloidal (zonal flows) structure formation processes [72]. Since zonal flows can be facilitated or locked in at the rational surface [103, 104], it might be also necessary to take the plasma current profile and its external constraints into account.
@@ Acknowledgments
This work was supported by: the Ministry of Science and Technology of the People’s Republic of China under Grant No. 2024YFE03190004, 2022YFE03100004; the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences under Award No. DE-FG02-04ER54738; the Sci DAC ABOUND Project, scw1832; the EPSRC under Grant No. EP/R014604/1; the National Natural Science Foundation of China under Grant Nos. 12375210 and 12305238; the Science and Technology Department of Sichuan Province under Nos. 2025ZNSFSC0059; the Southwestern Institute of Physics Project under No. 202301XWCX001-02; R&D Programs of ‘KSTAR Experimental Collaboration and Fusion Plasma Research (EN2501-16)’ and ‘High Performance Fusion Simulation R&D (EN2541-11)’ through Korea Institute of Fusion Energy (KFE) funded by the Government funds.

### v2

@@ 4.2 Models on the $\mathbf{E}\times\mathbf{B}$ staircase
Many theoretical models have been suggested to understand the physics of the $\mathbf{E}\times\mathbf{B}$ staircase. The $\mathbf{E}\times\mathbf{B}$ staircase was characterized as a mesoscopic transport state with co-existing super-diffusive transport events (avalanches) and permeable transport barriers (zonal flow layers) [68]. It was found in different frameworks [109, 106, 110] that such a state can be realized, though their physical interpretations may differ. In Garbet et al. [109], avalanches and zonal flow layers are the radially propagating wave packets and the wave trapping [111, 112, 113] quasi-periodic zonal flow solution in the system, respectively. In Milovanov et al. [106], zonal flow layers were described by coupled oscillators of nonlinear Schrödinger equation whose vibrations and sub-quadratic interaction allow a super-diffusive radial transport. In Chen et al. [110], drift wave solitons (akin to avalanches) can develop via the nonlinear drift wave and zonal flow interaction, and they are confined by spontaneously excited zonal flows.
@@
Some researches [86, 114, 87, 115, 69, 116, 117] have emphasized a role of mixing processes in the formation of mesoscopic zonal flows or profile corrugation. In Kosuga et al. [86], the pressure corrugation was understood as the jamming of avalanching heat flux due to the finite time delay [114]. In Ashourvan and Diamond [87] and Guo et al. [115], the bistability by dynamic mixing length was adopted to implement the inhomogeneous mixing and result in the density corrugation. In Wang et al. [69], the initial zonal flow driven by the Reynolds stress was fortified by the in-phase mean radial electric field variation during the temperature relaxation event. In Leconte and Kobayashi [116], the radial variation of the transport cross phase between the density and electric potential fluctuations (and the associated particle flux) can result in the density corrugation. In Yan and Diamond [117], alternating the enhanced and mitigated transport regions via the wave-particle resonance and non-resonance, respectively, due to the spatial structure of the mean flow lead to a corrugated profile in the trapped ion mode turbulence.
@@ 4.3 Some lessons from analogies
Revisiting details of specific models and comparing them with experimental results are left for future work. Here, in line with the theme of the workshop, “Layering—A structure formation mechanism in oceans, atmospheres, active fluids and plasmas”, the $\mathbf{E}\times\mathbf{B}$ staircase is viewed as one of self-organized layering phenomena in non-equilibrium complex systems and compared with other long-studied phenomena. Although there may be no universal laws for self-organized structures in complex systems, lessons can be learned by comparing observations in different systems [118].
@@
The $\mathbf{E}\times\mathbf{B}$ staircase is a pattern in the near-marginal regime characterized by the Kubo number ($Ku=\tau_{\mathrm{corr}}/\tau_{\mathrm{turn}}$, the ratio between the correlation time and the eddy turnover time) larger than 1. $Ku>1$ means more efficient eddy-particle interaction than decorrelation or detrapping processes, and results in the discrepancy from the local diffusive transport models [72]. Both the $\mathbf{E}\times\mathbf{B}$ staircase and the RB convection cells are patterns that form away from a state dictated by the homogeneous mixing of diffusive process (see Table 4). It demands us to go beyond the diffusive paradigm [36].
However, the $\mathbf{E}\times\mathbf{B}$ staircase is the secondary structure out of turbulent fluctuations, while the RB cells are more like steady laminar flows in the pattern of regular cells driven by the primary instability. While the scale of the RB convection cells is largely determined by the most unstable linear mode in the system, the nonlinear dynamics would matter for the scales of the $\mathbf{E}\times\mathbf{B}$ staircase, $\delta^{\mathrm{flow}}$ and $\Delta$. For example, $\delta^{\mathrm{flow}}$ can be extended by the turbulence spreading [120].
As the $\mathbf{E}\times\mathbf{B}$ staircase can be destroyed by avalanches (turbulent heat flux), the RB cells also become more complex, deformed or destructed, with turbulent fluctuations as the temperature difference is further increased (or, equivalently the viscosity is reduced). In Ramirez and Diamond [121], the effect of turbulent fluctuation on a staircase state of cellular flows was investigated, providing insight into the fluctuation strength condition that the staircase state can survive.
@@
Rayleigh-Bénard cells Zonal jets The $\mathbf{E}\times\mathbf{B}$ staircase
@@
Key condition $Ra=\frac{\tau_{\mathrm{cond}}}{\tau_{\mathrm{conv}}}>Ra_{\mathrm{crit}}$ $\tau_{\mathrm{wave}}\sim\tau_{\mathrm{turn}}$ [^4] Equating an eddy turnover time $\tau_{\mathrm{turn}}$ to the Rossby wave period $\tau_{\mathrm{wave}}$ yields a transition scale. Depending on how we define a turnover time, different scales can be found [122]. The Rhines scale [107] is obtained when $\tau_{\mathrm{turn}}$ is given by $\lambda_{\mathrm{corr}}/v$ where $\lambda_{\mathrm{corr}}$ and $v$ are the correlation length and typical velocity of the fluid eddy. $Ku=\frac{\tau_{\mathrm{corr}}}{\tau_{\mathrm{turn}}}>1$  [^5] Within a strict definition based on the original reference [67], the $\mathbf{E}\times\mathbf{B}$ staircase is not merely a set of simple shear flow layers, but rather shear flow layers that co-exist with avalanche transport. In other words, they are shear flow layers that originate and evolve in close relationships with avalanche transport. $Ku>1$ condition, indicative of the avalanche transport regime, would be a necessary but not sufficient condition for the $\mathbf{E}\times\mathbf{B}$ staircase. Note that this does not mean that zonal flows cannot exist in $Ku<1$ regime [123, 124].
@@
A more relevant and famous analogy has been made with zonal jets in the geophysical fluid [63, 125]. Indeed, the formal analogy can be made between the Charney equation for the (beta-plane) geophysical fluid dynamics and the Hasegawa-Mima equation for the (slab geometry) drift wave turbulence in plasmas.
In those simplified systems, there exists a conserved quantity called potential vorticity (PV; $Q$) attached to a fluid element, i.e. $\frac{dQ}{dt}\approx 0$ where $d/dt$ indicates the material derivative. For the total PV variation is preserved at zero within the system, if the PV mixing (by turbulent eddies or wave breaking) occurred in one location and reduce the PV gradient there, the PV gradient should be increased in other place. This is called the inhomogeneous PV mixing or PV staircase [66].
@@
One enlightenment from the analogy with the zonal jets is the non-acceleration theorem [127]. It states that steady and ideal (non-dissipative) waves or fluctuations have no net effect in changing the mean flow of the system, implying the important role of turbulence spreading flux (see item 2.2 of Section 2) in accelerating the zonal flow [103, 104]. For example, the turbulence influx through the plasma boundary can play a role to drive zonal flows in the core plasma (see item 2.4 of Section 2). Turbulence spreading can mediate the interaction between zonal flow layers [105] of the $\mathbf{E}\times\mathbf{B}$ staircase, suggesting an explanation for its long range order.
In realistic fusion plasmas, however, directly adopting the conventional picture of the inverse cascade is debatable. It relies on the large separation between the injection scale and the dissipation scale, but these scales may not be clearly distinguished in fusion plasmas with multi-scale instabilities. Also, there can be multiple processes generating a radial mixing zone in fusion plasmas [128, 129, 75], and defining the corresponding scale to the Rhines scale is not trivial. Avalanches are intrinsically the scale-free transport [71] unless they are limited by zonal flows. The width of the radial mixing region $\Delta$, the radial transport scale or the distance between the zonal flow layers, might be rather set by the competition or interaction between the radial (avalanches) and poloidal (zonal flows) structure formation processes [74]. Since zonal flows can be facilitated or locked in at the rational surface [90, 89, 91], it might be also necessary to take the plasma current profile and its external constraints into account.
@@ Acknowledgments
This work was supported by: the Ministry of Science and Technology of the People’s Republic of China under Grant No. 2024YFE03190004, 2022YFE03100004; the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences under Award No. DE-FG02-04ER54738; the Sci DAC ABOUND Project, scw1832; the EPSRC under Grant No. EP/R014604/1; the National Natural Science Foundation of China under Grant Nos. 12575237 and 12305238; the Nuclear Technology R&D Program under Grant No. HJSYF2024(02); the Science and Technology Department of Sichuan Province under No. 2025ZNSFSC0059; the Southwestern Institute of Physics Project under No. 202301XWCX001-02; R&D Programs of “High Performance Tokamak Plasma Research & Development (code No. EN2601-17)”, “Korea-US Collaboration Research for High Performance Plasma on Tungsten Divertor (EN2603-02)”, and “High Performance Fusion Simulation R&D (EN2641-12)” through the Korea Institute of Fusion Energy (KFE) funded by the Government funds, Republic of Korea; the National Research Foundation (NRF) of Korea under Grant No. RS-2023-00281272.